OpenStudy (anonymous):

Match each equation in the left column with its solution in the right column. Match Term Definition 3(x – 6) = 6 A) x = 2 3(2x – 1) = 15 B) x = 3 6(x + 1) = 12 C) x = 8 2(4x + 3) = 22 D) x = 1

OpenStudy (anonymous):

try plugging the values into x and mach them. it is very easy a cave man can do it :)

OpenStudy (texaschic101):

3(x - 6) = 6 -- distribute the 3 through the parenthesis 3x - 18 = 6 -- add 18 to both sides 3x = 6 + 18 -- combine like terms 3x = 24 -- divide by 3 x = 8 (C) 3(2x - 1) = 15 6x - 3 = 15 6x = 15 + 3 6x = 18 x = 3 (B) 6(x + 1) = 12 6x + 6 = 12 6x = 12 - 6 6x = 6 x = 1 (D) 2(4x + 3) = 22 8x + 6 = 22 8x = 22 - 6 8x = 16 x = 2 (A)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Nope Ty

OpenStudy (texaschic101):

your welcome :)

OpenStudy (anonymous):

It's diff from how we do it in Russia lol

OpenStudy (texaschic101):

really ? How do you do it there ?

OpenStudy (texaschic101):

By the way, you have a cool name. I love the name Anastasia

OpenStudy (anonymous):

All we do is divide then add and we get our answere I think we work backwards from u guyes

OpenStudy (anonymous):

Ty

OpenStudy (texaschic101):

glad I could help :)

OpenStudy (anonymous):

And Ty fore saying u like my name